The larger box is 8in. wide, 11 in. high, and 3 in. deep. How many square inches of cardboard are needed to make this box? Can you show me the formula?
Answer by checkley77(12569)

A box has 8 piecess (2 ends, 2 tops & 2 sides)
2*8*11=176^2 in.
2*8*3=48^2 in.
2*11*3=66^2 in.
176+48+66=290^2 in.